Suppose that the firm has two possible activities to produce output. Activity a uses a1 units of good 1 and a2 units of good 2 to produce 1 unit of output. Activity b uses bl units of good 1 and b2 units of good2 to produce 1 unit of output. Factors can only be used in these fixed proportions. If the factor prices are (wl, w2), what are the demands for the two factors? What is the cost function for this technology? For what factor prices is the cost function not differentiable?


1
Expert's answer
2021-02-10T15:32:25-0500
"Solution"

Here we derive the cost function with the help of the Cobb-Douglas production functions. Say, given "\alpha+\beta=1" corresponds to constant returns to scale, "\alpha+\beta <1" , showing decreasing returns and "\alpha+\beta>1", corresponds to increasing returns.

Now, suppose that "c(w,q)" is the cost function, then


"a_1(w_1,w_2,q)=q^\frac{1}{(\alpha+\beta)}(\frac{\alpha\ w_2}{\beta\ w_1})\\
a_2(w_1,w_2,q)=q^\frac{1}{\alpha+\beta}(\frac{\beta\ w_1}{\alpha\ w_2})\frac{\alpha}{(\alpha+\beta)}\\
c(w_1,w_2,q)=q\frac{1}{(\alpha+\beta)}[\frac{a}{\beta})\frac{\alpha}{(\alpha+\beta)}+(\frac{\alpha}{\beta})-\frac{\alpha}{\alpha+\beta}]w_1\frac{\alpha}{(\alpha+\beta)}w_2\frac{\beta}{\alpha+\beta}" is thus the cost function for this technology.


While "\theta\varnothing(w_1,w_2)" is the factor price for which the cost function is differentiable.